You are a bungee jumping fanatic and want to be the first bungee jumper on Jupiter. The length of your bungee cord is 45.0 m. Freefall acceleration on Jupiter is 23.1 m/s2. What is the ratio of your speed on Jupiter to your speed on Earth when you have dropped 45.0 m? Ignore the effects of air resistance and assume that you start at rest.
1
Expert's answer
2010-11-16T06:52:34-0500
Let's find the time of free fall on the Jupiter and the Earth:
t =√(2h/g1) = √(2*45/23.1) = 1.97 s
t =√(2h/g2) = √(2*45/9.8) = 3.03 sv(Jupiter) = 23.1*1.97 = 45.5 m/s
v(Earth) = 9.8*3.03 = 29.7 m/s
v(Jupiter) / v(Earth) = 45.5/29.7 = 1.5